Find the locus of the point whose distance from 

the point (0, 2) is five times its distance from the line x/3+y/4=1


1
Expert's answer
2021-01-13T19:34:06-0500

Let "(\alpha,\beta)" be the point whose distance from the point "(0,2)" is five times it's distance from the line "\frac {x}{3}+\frac{y}{4}=1" .

The distance between "(\alpha,\beta)" and "(0,2)" is "=\sqrt{(\alpha-0)^2+(\beta-2)^2}"

Given equation of the line "\frac {x}{3}+\frac{y}{4}=1" can be written as "4x+3y=12" ............(1)

Now distance from the point "(\alpha,\beta)" to the line is "=|\frac{4\alpha+3\beta-12}{\sqrt {(4)^2+(3)^2}}|" "=|\frac{4\alpha+3\beta-12}{5}|"

Therefore according to the problem, "\sqrt{(\alpha-0)^2+(\beta-2)^2}=5.|\frac{4\alpha+3\beta-12}{5}|"

"\implies \sqrt{(\alpha-0)^2+(\beta-2)^2}=|4\alpha+3\beta-12|"

"\implies (\alpha-0)^2+(\beta-2)^2=|4\alpha+3\beta-12|^2" [ Squaring both side]

"\implies \alpha^2+\beta^2-4\beta+4=16\alpha^2+9\beta^2+144+24\alpha\beta-96\alpha-72\beta"

"\implies 15\alpha^2+8\beta^2+24\alpha\beta-96\alpha-68\beta+140=0"

Hence required locus of the point "(\alpha,\beta)" is "15x^2+8y^2+24xy-96x-68y+140=0"

Which the required solution.